# Language Models and Logic Programs for Trustworthy Financial Reasoning

### v2

@@ Conclusion
We show how the integration of symbolic solvers with frontier language models can enhance the capability of AI tax assistance systems, and highlight their potential to improve equitable access to accurate and affordable tax guidance. Although tradeoffs remain between up-front costs of translating rules into formal logic versus ongoing inference-time computational and error costs, leveraging symbolic reasoning reduces overall expenses while improving auditability in both cases. Future research will explore how further scale or specialized smaller models optimized for faithful translation can improve aspects of this task, and develop methods for efficient translation of statutory rules into formal logic to enable effective few-shot learning. Finally, real-world implementation should be facilitated through interactive user studies using a further optimized system, enabling widespread accessibility to trusted, cost-efficient AI tax assistance.

### v3

@@ Conclusion
We frame statutory reasoning as a semantic parsing task, instructing language models to convert statutory and case text into executable code. We show how the integration of symbolic solvers with frontier language models can enhance the capability of AI tax assistance systems, and highlight their potential to improve access to accurate and affordable tax guidance. Although tradeoffs remain between up-front costs of translating rules into formal logic versus ongoing inference-time computational and error costs, leveraging symbolic reasoning reduces overall expenses while improving auditability in both cases. Future research will explore how further scale or specialized smaller models optimized for faithful translation can improve aspects of this task, and develop methods for efficient translation of statutory rules into formal logic to enable effective few-shot learning and help inform policy design.
Ethical Statement
Although we believe our results are promising, this paper should not be understood as a recommendation for individuals filing their taxes. Instead, we hope that policymakers and software developers will extend these insights to reduce the individual effort required for tax filing, building systems which can allow a single human-in-the-loop to complete filings more efficiently. Such steps would require careful consideration of liability issues, impacts on vulnerable populations, and non-financial costs associated with tax reviews such as stress or time. Furthermore, the reliance on closed-source models may place private companies at a choke point, so further advances in the open-source domain are essential.
Acknowledgments
This work was funded in part by the U.S. National Science Foundation under grant No. 2204926, and by the Defense Advanced Research Project Agency (DARPA) CODORD program. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or DARPA.